If the width of a rectangle is 5 inches and its perimeter is 40 inches, find the
length of the rectangle.

Answer by stanbon(75887)

If the width of a rectangle is 5 inches and its perimeter is 40 inches, find the
length of the rectangle.
------------------
P = 2(L+W)
40 = 2(L+5)
20 = L+5
Length = 15 inches
======================
Cheers,
Stan H.
======================